OpenStudy (anonymous):

Solve the polynomial equation. 16n2 - 25 = 0

OpenStudy (anonymous):

add 25 divide by 16 take the square root (don't forget the \(\pm\))

OpenStudy (anonymous):

n=+- 5/4? @satellite73

OpenStudy (nirmalnema):

16n^2-25=0 16n^2=25 n^2=25/16 n=\[\pm \sqrt{25/16}\] n=+-5/4